# Accelerating Antibiotic Discovery with Large Language Models and Knowledge Graphs

### v1

@@ 4.1 Natural products literature: descriptive bibliometric analysis
Assessing the size and growth of the natural products and antibiotics literature is crucial to highlight the extensive effort required by reviewers. In 2024, it is more than 50,000 new articles that have been indexed in PubMed for the searches "natural products" and "antibiotics", reporting novel links between organisms, chemicals, and activities. While keeping up with new literature is crucial, Figure 4.A shows that a significant portion of annotated relations in the LOTUS database comes from older articles (pre-1970). Given the evolution of taxonomy and nomenclature over time, relying on original organism identifications from the text is unreliable, making synonym resolution essential for linking past and novel relations. Using the publicly available literature from PubMed as a reference for an alert system also requires evaluating its coverage. Although PubMed includes over 38 millions articles, Figure 4.B indicates that fewer than half of the annotated references in the LOTUS database are actually indexed in PubMed. This observation underscores a notable gap in PubMed’s coverage. Nevertheless, given the extensive volume of literature within PubMed, it’s also reasonable to expect that many relevant references may be missing from LOTUS. Also, while we observed that most articles are in English, this likely reflect a bias from the resources used in LOTUS, and, other corpora) are also expected to be relevant.
@@ 4.2 Pipeline execution
S/ strictum * Cytosporone E 29354097 22690142 ✓/× Strong
@@
H. aurantius Hypomycetin ext. ref(2) ext. ref(2) ×/× Missed
@@
P. byssoides Pericosine A 18043803 26928999 ✓/× Strong
@@ 4.3 Evaluation on Discarded Hits
Among the 73 initial identifications, 12 were discarded as negative hits after an extensive manual review. Figure 5 displays the distribution of alerts raised for each discarded organism from CL (left) and OL (center) evidence. While the number of alerts varies (max: 174, min: 1), each organism has at least one Strong alert. To assess the impact of the extraction pipeline, an ablation study (Figure 5 right) using only LOTUS database annotations showed that only 5 of the 12 negative hits could be identified, highlighting the added value of the RE step. For the 12 negative hits, the reviewers previously identified 27 evidence triples (organism-chemical-activity). Table 1 compares these with system-generated alerts from Figure 5, focusing on chemical-based alerts, as all evidence provided by the reviewers are linked to a chemical. An alert is considered missed if the chemical was not retrieved (via RE or LOTUS) or its activity was not reported [^5] Neither Strong, Medium. Among the 27 reviewer-reported evidence, 6 were missed by the system, including 3 because of non-indexed references or unavailable texts in PubMed. Notably, 25 of the 27 chemicals were successfully retrieved, with 22 through the RE step. A detailed error analysis is provided in Appendix D. Except for Acremonium butyri, all negative hits were correctly discarded. Screenshots of the user interface, including an example for Sarocladium strictum, are shown in Appendix E.
@@ Appendix D Error analysis
Hypomyces aurantius - Hypomycetin: The reference article found by the reviewers (DOI:10.3891/acta.chem.scand.51-0855) is not indexed in PubMed, nor LOTUS.

### v2

@@ 4.1 Natural products literature: descriptive bibliometric analysis
Assessing the size and growth of the natural products and antibiotics literature is crucial to highlight the extensive effort required by reviewers. In 2024, it is more than 50,000 new articles that have been indexed in PubMed for the searches "natural products" and "antibiotics", reporting novel links between organisms, chemicals, and activities. While keeping up with new literature is crucial, Figure 4.A shows that a significant portion of annotated relations in the LOTUS database comes from older articles (pre-1970). Given the evolution of taxonomy and nomenclature over time, relying on original organism identifications from the text is unreliable, making synonym resolution essential for linking past and novel relations. Using the publicly available literature from PubMed as a reference for an alert system also requires evaluating its coverage. Although PubMed includes over 38 millions articles, Figure 4.B indicates that fewer than half of the annotated references in the LOTUS database are actually indexed in PubMed. This observation underscores a notable gap in PubMed’s coverage. Nevertheless, given the extensive volume of literature within PubMed, it’s also reasonable to expect that many relevant references may be missing from LOTUS. Also, while we observed that most articles are in English, this likely reflect a bias from the resources used in LOTUS, and, other (eg. traditional Chinese medicine prescriptions) corpora are also expected to be relevant.
@@ 4.2 Pipeline execution
S. strictum * Cytosporone E 29354097 22690142 ✓/× Strong
@@
H. aurantius Hypomycetin ext. ref(2) ext. ref(2) ×/✓ Missed
@@
P. byssoides Pericosine A 18043803 26928999 ✓/✓ Strong
@@ 4.3 Evaluation on Discarded Hits
Among the 73 initial identifications, 12 were discarded as negative hits after an extensive manual review. Figure 5 displays the distribution of alerts raised for each discarded organism from CL (left) and OL (center) evidence. While the number of alerts varies (max: 174, min: 1), each organism has at least one Strong alert. To assess the impact of the extraction pipeline, an ablation study (Figure 5 right) using only LOTUS database annotations showed that only 5 of the 12 negative hits could be identified, highlighting the added value of the RE step. For the 12 negative hits, the reviewers previously identified 27 evidence triples (organism-chemical-activity). Table 1 compares these with system-generated alerts from Figure 5, focusing on chemical-based alerts, as all evidence provided by the reviewers are linked to a chemical. An alert is considered missed if the chemical was not retrieved (via RE or LOTUS) or its activity was not reported [^5] Neither Strong, Medium. Among the 27 reviewer-reported evidence, 6 were missed by the system, including 3 because of non-indexed references or unavailable texts in PubMed. Notably, 26 of the 27 chemicals were successfully retrieved, with 22 through the RE step. A detailed error analysis is provided in Appendix D. Except for Acremonium butyri, all negative hits were correctly discarded. Screenshots of the user interface, including an example for Sarocladium strictum, are shown in Appendix E.
@@ Appendix D Error analysis
Hypomyces aurantius - Hypomycetin: The reference article identified by the reviewers (DOI:10.3891/acta.chem.scand.51-0855) is indexed in LOTUS. This article also describes the antifungal activity of Hypomycetin. However, since the article is not indexed in PubMed, the evidence of its activity has not been extracted.